OpenStudy (anonymous):

Write an equation of the line through (3,-2) and perpendicular to 5y=x-10.

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

do you have an idea how to do it

OpenStudy (anonymous):

I'm actually having a tough time with linear inequalities and graphing

OpenStudy (anonymous):

slope = 1/5 therefore (y+2)/(x-3)=1/5 now solve it and get the answer '

OpenStudy (anonymous):

sorry guys slope is -5 else everything is same

OpenStudy (anonymous):

The problem is, is that I don't understand how to solve it or how you ended up with -5.

OpenStudy (anonymous):

since line is perpendicular to the given line therefore the product of two slopes is minus one therefore if given line has a slope 1/5 other which we are going to find that must be -5

OpenStudy (anonymous):

Yes. Slope of perpendicular line is - 1/slope = -5. Next solve y = -5x + b for b, knowing x=3, y=-2

OpenStudy (anonymous):

So it would be -2 = -5x + 3? Did I get that correct?

OpenStudy (anonymous):

or wait. I meant -2=-5x(3) + b

OpenStudy (anonymous):

5y = x -10 can be rewritten as y=1/5 x -2. the slope of this equation is 1/5. The slope of a line perpendicular would be (i think it's called the recpriocal) -5. now use the equation: (Y-y) = m(X-x) where y=-2 and x=3.

OpenStudy (anonymous):

yeah very right

OpenStudy (anonymous):

y-2=-5(x-3)?

OpenStudy (anonymous):

close. look at y-2.

OpenStudy (anonymous):

y-(-2)=-5(x-3)

OpenStudy (anonymous):

or y+2=-5(x-3) which I think would then come out to: y+2=-5x+15

OpenStudy (anonymous):

In high school, my teachers wanted it written: y = -5(x-3)-2, because that is how it is easiest to "see" the graph.

OpenStudy (anonymous):

Thank you to everyone who has helped. I really appreciate you all for taking the time and effort to give me insight into this problem.